OpenStudy (dmezzullo):

]Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9)

OpenStudy (dmezzullo):

@jazy

OpenStudy (anonymous):

slope = 2; y intercept at (0, -4)

OpenStudy (dmezzullo):

tyvm but could u explain tht?

OpenStudy (anonymous):

\[6x-3y=12\]\[\frac{ -3y }{ -3 }= \frac{ -6x }{ -3 } +\frac{ 12 }{ -3 }\] \[y=2x-4\] Slope= 2 Y- intercept= 0,-4